# ClavaDDPM: Multi-relational Data Synthesis with Cluster-guided Diffusion Models

### v1

@@ B.2 Diagram for two-table relational databases
B.3 End-to-end algorithms for more tables
@@
4:     $C_{i,j}\leftarrow Clustering\left(X;Y\right)$triangleright Relation-aware clustering in Section 4.3.1
@@
Algorithm 3 takes in learned diffusion models $\mathcal{D}$, classifiers $\mathcal{C}$, group size distributions $\mathcal{S}$, and the DAG representation of the database $\mathcal{G}$, and outputs the synthetic database ~=$\tilde{\mathcal{R}}=\left\{\tilde{R}_{1},\ldots,\tilde{R}_{m}\right\}$. We first initialize the synthetic augmented tables to be empty (line 1). Then, for root augmented tables, since they have no parents to condition on, they can be directly synthesized from their diffusion models (line 2-4). Next, we traverse the database in topological order to synthesize the remaining augmented tables (line 5-16): If we have already synthesized $\tilde{T}_{i}$ before, which means we encounter the multi-parent dilemma, we just store the old version and continue to generate a new version (line 6-9). For each parent-child relationship $R_{i}\rightarrow R_{j}$, we must have already sampled the augmented parent table $\tilde{T}_{j}$. This is because we follow the topological order of a DAG, and all root augmented tables have been synthesized as base cases. Therefore, we can obtain the synthetic latent variables $\tilde{C}_{i,j}$ from the synthetic augmented parent $\tilde{T}_{j}$ (line 10). Then, we iterate through each synthetic latent value $\tilde{c}_{i,j}$ and perform a two-step sampling: (1) use the learned group size distribution to conditionally sample a group size $\tilde{s}$ (line 12); (2) sample $\tilde{s}$ rows of data conditioned on $\tilde{c}_{i,j}$ using classifier guided sampling (line 13). We repeat this process until the augmented child table $\tilde{T}_{i}$ is fully synthesized. We simply obtain synthetic tables from synthetic augmented tables by removing all synthetic latent columns (line 17-19). Finally, for all the encountered multi-parent dilemmas, we follow 4.3.3 to match the divergent versions.
@@ C.3.2 Hyper Parameters
We list the major hyper parameters used by ClavaDDPM for each dataset in  5, and we provide an empirical guidance for hyper parameter tuning: it is suggested to use number of clusters $k$ to be at least $20$, and classifier scale $\eta$ to be in $\left[0.5,2\right]$. We empirically find ClavaDDPM consistently perform well in such a range across all datasets. Parent scale $\lambda$ is a less sensitive factor, and $\lambda=1$ is a stable starting point for tuning. In general, ClavaDDPM is robust, with a small hyper parameter space, and there is very little need for tuning.
@@ D.1 Agree rate discussion
As introduced in Section 4.3.1, within each foreign key group, we perform a majority voting to synchronize the assigned cluster label among the group. To measure the consistency of such majority voting process, we introduce the measurement of agree rate, which computes the average ratio of agreeing on the mode within each foreign key group, and the metric avg agree-rate is the average of all per-table agree rates withing a multi-table dataset. However, our experiment results in Table 2 indicate that the relation-aware clustering process is robust against such factors, and the GMM model is capable of capturing nuances in data patterns. In our experiments on Berka dataset, ClavaDDPM’s clustering process achieves a consistent agree rate around $45\%$, which is practically high enough given we have $20$ clusters. Intuitively, when parent scale approaches infinity, the clustering is performed completely on parent table, which will lead to a perfect agree rate. Also note that a higher agree rate does not always imply a better performance, and the disagreement can potentially come from the intrinsic parent-child relationships. e.g. when child data is intrinsically independent of parent data, it is reasonable to have noisy learned latent variables, leading to low agree rates. However, in such cases the noisy latent variable would not degenerate model performance, because the best strategy will be a direct sampling of child table, rather than conditioning on some enforced prior distribution. In addition, as shown in 2, agree rates are highly affected by the number of clusters chosen, and there exists a trade-off between the granularity in clustering and consistency. In an extreme case where we have $k=1$ cluster, indicating an infinitely coarse-grained latent learning, it trivially achieves perfect agree rates.
D.2 High-order Single-table Evaluation
@@
We evaluated high-order single-table metrics on the California dataset across all baseline models and ClavaDDPM. Following [50], for the evaluation of MLE we perform a $90\%,10\%$ train-test split, where the $F_{1}$ and $R^{2}$ metrics are evaluated on the $10\%$ holdout set. Note that although PrivLava has an advantage on the California dataset when evaluated with statistical tests  1, ClavaDDPM exhibits competitive, or even stronger performance than PrivLava on higher-order metrics. Especially for MLE, ClavaDDPM surpasses PrivLava by $13.07\%$ in terms of average $R^{2}$ in Individual table, and also beats PrivLava on average $F_{1}$ in both tables. Also notice that the baseline ST-ClavaDDPM dominates in high-order metric evaluations, demonstrating the strength of our Gaussian diffusion-only backbone model.
@@
D.3 Privacy Sanity Check
@@
As shown in table 6, ClavaDDPM although neither specialized in privacy preserving, nor in single table synthesis, it still maintains a reasonable privacy level. The charts 3 demonstrates the distributions of DCR scores, where ClavaDDPM is in blue. The overall distribution is more leaning to the right side, indicating an overall higher DCR distribution.

### v2

@@ B.1 Diagram for two-table relational databases
B.2 End-to-end algorithms for more tables
@@
4:     $C_{i,j}\leftarrow Clustering\left(X;Y\right)$triangleright Relationship-aware clustering in Section 4.3.1
@@
Algorithm 3 takes in learned diffusion models $\mathcal{D}$, classifiers $\mathcal{C}$, group size distributions $\mathcal{S}$, and the DAG representation of the database $\mathcal{G}$, and outputs the synthetic database ~=$\tilde{\mathcal{R}}=\left\{\tilde{R}_{1},\ldots,\tilde{R}_{m}\right\}$. We first initialize the synthetic augmented tables to be empty (line 1). Then, for root augmented tables, since they have no parents to condition on, they can be directly synthesized from their diffusion models (line 2-4). Next, we traverse the database in topological order to synthesize the remaining augmented tables (line 5-16): If we have already synthesized $\tilde{T}_{i}$ before, which means we encounter the multi-parent dilemma, we just store the old version and continue to generate a new version (line 6-9). For each parent-child relationship $R_{i}\rightarrow R_{j}$, we must have already sampled the augmented parent table $\tilde{T}_{j}$. This is because we follow the topological order of a DAG, and all root augmented tables have been synthesized as base cases. Therefore, we can obtain the synthetic latent variables $\tilde{C}_{i,j}$ from the synthetic augmented parent $\tilde{T}_{j}$ (line 10). Then, we iterate through each synthetic latent value $\tilde{c}_{i,j}$ and perform a two-step sampling: (1) use the learned group size distribution to conditionally sample a group size $\tilde{s}$ (line 12); (2) sample $\tilde{s}$ rows of data conditioned on $\tilde{c}_{i,j}$ using classifier guided sampling (line 13). We repeat this process until the augmented child table $\tilde{T}_{i}$ is fully synthesized. We simply obtain synthetic tables from synthetic augmented tables by removing all synthetic latent columns (line 17-19). Finally, for all the encountered multi-parent dilemmas, we follow Section 4.3.3 to match the divergent versions.
@@ C.3.2 Hyper Parameters
We list the major hyper parameters used by ClavaDDPM for each dataset in Table  5, and we provide an empirical guidance for hyper parameter tuning: it is suggested to use number of clusters $k$ to be at least $20$, and classifier scale $\eta$ to be in $\left[0.5,2\right]$. We empirically find ClavaDDPM consistently perform well in such a range across all datasets. Parent scale $\lambda$ is a less sensitive factor, and $\lambda=1$ is a stable starting point for tuning. In general, ClavaDDPM is robust, with a small hyper parameter space, and there is very little need for tuning.
@@ D.1 Agree rate discussion
As introduced in Section 4.3.1, within each foreign key group, we perform a majority voting to synchronize the assigned cluster label among the group. To measure the consistency of such majority voting process, we introduce the measurement of agree rate, which computes the average ratio of agreeing on the mode within each foreign key group, and the metric avg agree-rate is the average of all per-table agree rates within a multi-table dataset.
$A(g)=\frac{m_{g}}{|g|}$ (11)
$A(g)$ represents the agree rate of some group $g$, where $m_{g}$ represents the number of records that are assigned the mode cluster within $g$, and the avg agree-rate
$A_{\text{AVG}}=\frac{1}{|G|}\sum_{g\in G}A(g)$ (12)
is computed as the average of the agree rates across all groups.
However, our experiment results in Table 2 indicate that the relationship-aware clustering process is robust against such factors, and the GMM model is capable of capturing nuances in data patterns. In our experiments on Berka dataset, ClavaDDPM’s clustering process achieves a consistent agree rate around $81\%$, which is practically high enough given we have $20$ clusters. Intuitively, when parent scale approaches infinity, the clustering is performed completely on parent table, which will lead to a perfect agree rate. Also note that a higher agree rate does not always imply a better performance, and the disagreement can potentially come from the intrinsic parent-child relationships. e.g. when child data is intrinsically independent of parent data, it is reasonable to have noisy learned latent variables, leading to low agree rates. However, in such cases the noisy latent variable would not degenerate model performance, because the best strategy will be a direct sampling of child table, rather than conditioning on some enforced prior distribution. In addition, as shown in Table 2, agree rates are highly affected by the number of clusters chosen, and there exists a trade-off between the granularity in clustering and consistency. In an extreme case where we have $k=1$ cluster, indicating an infinitely coarse-grained latent learning, it trivially achieves perfect agree rates.
D.2 Selecting Number of Clusters $k$
We conducted finer-grained experiments to examine the effect of the number of clusters $k$ on model performance, as shown in Figure 3, which offer empirical insights for selecting $k$. Based on the results, (1) a binary search approach could be used to efficiently find a suitable $k$, and (2) while it may require more computational resources, opting for a larger $k$ is generally a safe choice.
Figure 3: Smoothed model performance on Berka dataset regarding different $k$ (measured by AVG 2-way), where $k=\infty$ represents assigning each row a unique class.
D.3 High-order Single-table Evaluation
@@
We evaluated high-order single-table metrics on the California dataset across all baseline models and ClavaDDPM. Following [50], for the evaluation of MLE we perform a $90\%,10\%$ train-test split, where the $F_{1}$ and $R^{2}$ metrics are evaluated on the $10\%$ holdout set. Note that although PrivLava has an advantage on the California dataset when evaluated with statistical tests (Table  1), ClavaDDPM exhibits competitive, or even stronger performance than PrivLava on higher-order metrics. Especially for MLE, ClavaDDPM surpasses PrivLava by $13.07\%$ in terms of average $R^{2}$ in Individual table, and also beats PrivLava on average $F_{1}$ in both tables. Also notice that the baseline ST-ClavaDDPM dominates in high-order metric evaluations, demonstrating the strength of our Gaussian diffusion-only backbone model.
@@
Table 6: High-level single-table metrics evaluated on the Household table and the Individual table from the California dataset.
D.4 Privacy Sanity Check
@@
As shown in table 7, ClavaDDPM although neither specialized in privacy preserving, nor in single table synthesis, it still maintains a reasonable privacy level. The charts 4 demonstrates the distributions of DCR scores, where ClavaDDPM is in blue. The overall distribution is more leaning to the right side, indicating an overall higher DCR distribution.
@@
Appendix E Complexity Analysis
Given a multi-relational database $\mathcal{G}=(\mathcal{R},\mathcal{E})$, with $m$ tables, $n$ foreign key constraints, and $p$ rows per table. For a $p$-row table, we denote the time complexity of performing GMM clustering as $c_{\text{GMM}}(p)$, training a diffusion model as $c_{\text{diff}}(p)$, training a classifier as $c_{\text{class}}(p)$, synthesizing as $c_{\text{syn}}(p)$, ANN searching as $c_{\text{ANN}}(p)$.
Phase 1: latent learning and table augmentation
$n\cdot c_{\text{GMM}}(p)$ (13)
Phase 2: training
$n\cdot c_{\text{class}}(p)+m\cdot c_{\text{diff}}(p)$ (14)
Note that in practice this phase is dominated by diffusion training, primarily influenced by $m$.
Phase 3: synthesis
$n\cdot c_{\text{syn}}(p)$ (15)
Additional step: matching
$n\cdot c_{\text{ANN}}(p)$ (16)
Note that the runtime in this phase is negligible compared to the earlier phases, particularly with the FAISS implementation in the non-unique matching setup.
Total
$n\left(c_{\text{GMM}(p)+c_{\text{class}}}(p)+c_{\text{syn}}(p)+c_{\text{ANN}}(p)\right)+mc_{\text{diff}}(p)$ (17)
the overall runtime is dominated by Phase 2 (training) and Phase 3 (synthesis), with the critical factors being $m$, $n$, and $p$. The model remains robust against the number of clusters in Phase 1, as the impact on runtime is minimal due to the dominance of the later phases.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We provide empirical evidence in the evaluation section for the claims made in the abstract and introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: In the conclusion section (second paragraph), we discuss the limitations of our work and our interest in addressing them as future work.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We list all the assumptions required to derive our generative process. We do not have any proof.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We detail the datasets, baseline algorithms, implementation details of our algorithms, and evaluation metrics in the appendix.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material? [Yes] Justification: We upload supplementary materials including code for reproducibility.
  23. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  24. 6.
    Experimental Setting/Details
  25. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  26. Answer: [Yes]
  27. Justification: We provide the basic experimental setting/details, ablation study in the evaluation section, and details in the appendix.
  28. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  29. 7.
    Experiment Statistical Significance
  30. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  31. Answer: [Yes]
  32. Justification: We repeat experiments for each configuration three times and report their averaged performance with standard deviations.
  33. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  34. 8.
    Experiments Compute Resources
  35. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  36. Answer: [Yes]
  37. Justification: We include computing specifics in the evaluation section.
  38. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  39. 9.
    Code Of Ethics
  40. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethic https://neurips.cc/public/EthicsGuidelines?
  41. Answer: [Yes]
  42. Justification: We carefully go through the NeurIPS Code of Ethic and ensure we follow it.
  43. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  44. 10.
    Broader Impacts
  45. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  46. Answer: [Yes]
  47. Justification: We discuss the broader impacts of our work in the conclusion section.
  48. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  49. 11.
    Safeguards
  50. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  51. Answer: [N/A]
  52. Justification: We use public datasets for experiments.
  53. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  54. 12.
    Licenses for existing assets
  55. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  56. Answer: [Yes]
  57. Justification: We add citations and resources for the datasets and source code used in the evaluation.
  58. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  59. 13.
    New Assets
  60. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  61. Answer: [Yes]
  62. Justification: We document our new code and and provide references for all code adopted from public resources.
  63. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  64. 14.
    Crowdsourcing and Research with Human Subjects
  65. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  66. Answer: [N/A]
  67. Justification: Our paper does not involve crowdsourcing nor research with human subjects.
  68. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  69. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  70. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  71. Answer: [Yes]
  72. Justification: Our paper does not involve crowdsourcing nor research with human subjects.
  73. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.